# A phase-space view of vibrational energies without the Born-Oppenheimer framework

### v1

@@ III Demonstration through a model problem
  1. 1.
    Exact calculations, where we diagonalize $\hat{H}$ over a joint grid in the nuclear coordinate $R$ and the electronic coordinate $r$.
  2. 2.
    BO calculations, where we loop over a grid of nuclear positions $R_{i}$, use a grid of electronic positions to diagonalize the electronic Hamiltonian associated to each nuclear grid point $\hat{H}_{\rm el}(R_{i})\ket{\psi_{0}}=\epsilon_{0}(R_{i})\ket{\psi_{0}}$, and then diagonalize the nuclear Hamiltonian $\hat{P}^{2}/2M+\epsilon_{0}(\hat{R})$ over the nuclear grid.
  3. 3.
    PS calculations, where we loop over a joint grid of nuclear positions and momenta $(R_{i},P_{j})$, use a grid of electronic positions to diagonalize the phase-space Hamiltonian (Eq. 32) associated to each joint grid point $\hat{H}_{W}^{\rm PS}(R_{i},P_{j})\ket{\psi_{0}}=\epsilon^{PS}_{W,0}(R_{i},P_{j})\ket{\psi_{0}}$, invert $\epsilon^{PS}_{W,0}(R_{i},P_{j})$ to a function in real space $\epsilon_{0}^{\rm PS}(R_{i},R_{j})$ using Eq. 28, and then diagonalize $\epsilon_{0}^{PS}(R_{i},R_{j})$ over the nuclear grid.
Below, we will also report results where we consider harmonic BO and PS surfaces, i.e. where we force $\epsilon_{0}(R)$ or $\epsilon^{\rm PS}_{W,0}(R,P)$ to be a quadratic function. More numerical details for these simulations can be found in SI.
@@
To begin our analysis, in Fig. 1(b), we plot the ground BO potential energy surface that arises from treating the light and heavy degrees of freedom (DoF) separately, i.e., treating the oxygen atoms as nuclei and the hydrogen atom as an electron. One should not expect this BO separation will not be a good approximation if the mass ratio $M/m$ is not large enough, such as the actual masses of oxygen and hydrogen atoms (16 and 1 atomic mass units, respectively). Thus, by varying the nuclei/electron mass ratio $M/m$, we can control non-adiabatic effects for this model Hamiltonian.
@@
In Fig. 2, we report our results for the ground to first excited vibrational energy gap as a function of the different mass ratio. For all methods, the absolute values of energy gap decreases as the mass ratio increases and converge to the same results in the large mass ratio limit ($\sim$1000 corresponding to a typical nuclear/electronic mass ratio) in Fig. 2(a). Thus, in Fig. 2(b), we plot the error of the energy gap as compared to the exact results. In the small mass ratio limit ($\sim$10 corresponding to the oxygen and hydrogen atoms), one can clearly identify that the phase-space method shows a much smaller error compared to the BO method. Notably, even within a harmonic approximation, the phase-space method still outperforms the BO and sometimes can be comparable to the quantum BO results. Lastly, to best understand these results, the inset of Fig. 2(b) shows the relative errors in the logarithm scale. As one might expect, the phase-space and BO energy errors both scale with the same power law ($\sim m/M$), but the phase-space method consistently gives an order of magnitude smaller relative error even in the large mass ration limit.
@@
It must be emphasized that, with a phase-space Hamiltonian, one can extract other useful properties as well (besides energy) and these quantities are also more accurate than the corresponding BO Hamiltonian. For example, in Fig. 3, we show the errors of expectation values of a few nuclear and electronic observables. First, for the electronic momentum, it is important to note that BO calculations always predict zero, as $\bra{\psi_{k}}\hat{\bm{p}}_{\rm e}\ket{\psi_{k}}=0$ for all BO electronic states at any nuclear geometry. However, as shown in Fig. 3(a), the phase-space method can capture the electronic momentum between the ground and first excited vibrational states with quantitative accuracy. Second, for nuclear momentum, Fig. 3(b) demonstrates that phase-space results are also much more accurate than BO methods. One interesting nuance is that one must replace the nuclear momentum operator $\hat{\bm{P}}$ by $\hat{\bm{P}}-\hat{\bm{\Gamma}}$ in the phase-space formalism which represents a canonical – kinetic momentum correspondence. Overall, for the off-diagonal electronic and nuclear momentum (between the same nuclear vibrational eigenstates), the phase-space method performs very well even with the harmonic approximation. As a side note, the on-diagonal nuclear and electronic momentum always vanish due to time reversal symmetry of the whole system (nuclei + electron). Finally, in Fig. 3(c) and (d), while the nuclear and electronic position observables predicted by BO and phase-space methods are not very different, the phase-space methods do provide slightly better results for the excited vibrational states. For further analysis, and a decomposition of the importance of different terms in the phase-space Hamiltonian, see the SI.
@@
IV Discussion and Conclusions
@@
Looking forward, it would appear clear from the plethora of data collected here and in Refs.22, 23 that a phase-space approach is more accurate than a BO approach – and with roughly the same cost. That being said, one must wonder: in what circumstances will the phase-space approach yield meaningfully different answers than BO theory? To that end, one can imagine several exciting future problems where one would like to go beyond the limitation of BO theory. The first potential arena would be to study proton transfer[34] and proton coupled electron/energy transfer[35, 36] where one could combine a phase-space approach with electronic structure methods that treat both light nuclei (protons) and electrons quantum mechanically (e.g., multi-component methods[37, 38, 39] and nuclear-electronic orbital methods (NEO)[40, 41]. Second, it will be fruitful to apply a phase-space approach to study magnetic molecules and solids, where spin-symmetry breaking can lead to intrinsic chiral phonon modes and alter vibrational frequencies[42, 43], and electronic momentum is known to be critically important in such circumstances[44, 45]. Third, given the ab initio and semiclassical nature of the phase-space electronic Hamiltonian, one immediately goal is to investigate nuclear (vibrational) effects at curve crossings during electron transfer processes[46] and decipher whether phase-space approaches (that include angular momentum conservation[47]) can help explain the nature of nuclear effects in chiral-induced spin selectivity[48, 49, 50]. Fourth, the present phase-space approach opens new and inexpensive avenues for exploring electron-phonon interactions, which could lead to a better atomistic understanding of superconductivity[51, 52, 53]. At the end of the day, there is a very real chance that phase-space electronic Hamiltonian approaches are here to stay and that a new world of very novel chemistry awaits us.

### v2

@@ 5.2 Methods
5.2.1 Exact calculations
For the exact results, we diagonalize $\hat{H}$ over a joint $(N_{r}\times N_{R})\times(N_{r}\times N_{R})$ grid in the nuclear coordinate $R$ and the electronic coordinate $r$ , where we choose $N_{r}=400$ grids points for the electronic coordinate $r\in\left[-2,2\right]\textup{\AA}$ and $N_{R}=100$ grid points for nuclear coordinate $R\in\left[2,4\right]\textup{\AA}$. Here and below (for all methods), we constructed the electronic momentum and electronic kinetic energy operators by the finite-difference stencil technique, while the nuclear kinetic operator is built with a Fourier transform using the conjugate momentum grid points.
5.2.2 BO calculations
For the BO results, we use the same grids as above the for the exact calculations (with $N_{r}=400$ grids points for the electronic coordinate $r\in\left[-2,2\right]\textup{\AA}$ and $N_{R}=100$ grid points for nuclear coordinate $R\in\left[2,4\right]\textup{\AA}$). As is standard, we begin by looping over the grid of nuclear positions $R_{i}$; thereafter we use the grid of electronic positions to diagonalize the electronic Hamiltonian associated for each nuclear grid point $\hat{H}_{\rm el}(R_{i})\ket{\psi_{0}}=E_{0}(R_{i})\ket{\psi_{0}}$. Finally, we diagonalize the nuclear Hamiltonian $\frac{\hat{P}^{2}}{2M}+E_{0}(\hat{R})$ over the nuclear grid to generate vibrational energies.
5.2.3 Phase-space Calculations
To generate the phase-space results, we first construct the phase-space electronic Hamiltonian for the three-particle model system in the MCM coordinates. Since the model Hamiltonian only has one-spatial dimension, we do not need to consider the rotational part of $\hat{\bm{\Gamma}}$ operator ($\hat{\bm{\Gamma}}^{\prime\prime}$). In the original coordinates, the phase-space Hamiltonian can be expressed as:
$\hat{H}_{\rm PS}(R_{1},R_{2},P_{1},P_{2})=\frac{(P_{1}-i\hbar\hat{\Gamma}_{1})^{2}}{2M}+\frac{(P_{2}-i\hbar\hat{\Gamma}_{2})^{2}}{2M}+\frac{\hat{p}^{2}}{2m}+\hat{V}(R_{1},R_{2},\hat{r}_{\rm e}).$ (73)
At this point, it will be convenient to define the nuclear center of mass (NCM) coordinates:
$\hat{R}_{\rm NCM}=\frac{\hat{R}_{1}+\hat{R}_{2}}{2}.$ (74)
If we change coordinates from $(R_{1},R_{2})$ to $(R,R_{NCM})$, the result is:
$\hat{H}_{\rm PS}(R,P,P_{NCM})=\frac{(P_{NCM}+\hat{p})^{2}}{2(M+M)}+\frac{\left(P-i\hbar\frac{(\hat{\Gamma}_{1}-\hat{\Gamma}_{2})}{2}\right)^{2}}{2\mu}+\frac{\hat{p}^{2}}{2m}+\hat{V}(R,\hat{r}).$ (75)
We can identify the $\hat{\Gamma}$ operator for the relative $R$ coordinate,
$\hat{\Gamma}=\frac{\hat{\Gamma}_{1}-\hat{\Gamma}_{2}}{2}.$ (76)
At this point, we ignore the MCM motion encapsulated in the first term in Eq. 75, so that the final phase-space electronic Hamiltonian becomes
$\hat{H}_{\rm PS}(R,P)=\frac{(P-i\hbar\hat{\Gamma})^{2}}{2\mu}+\frac{\hat{p}^{2}}{2m}+\hat{V}(R,\hat{r}),$ (77)
where
$\hat{\Gamma}_{1}=\frac{1}{2i\hbar}\left(\hat{\theta}_{1}\hat{p}+\hat{p}\hat{\theta}_{1}\right),\quad\hat{\Gamma}_{2}=\frac{1}{2i\hbar}\left(\hat{\theta}_{2}\hat{p}+\hat{p}\hat{\theta}_{2}\right)$ (78)
with
$\hat{\theta}_{1}=\frac{e^{-|\hat{r}-\frac{R}{2}|^{2}/\sigma^{2}}}{e^{-|\hat{r}-\frac{R}{2}|^{2}/\sigma^{2}}+{e^{-|\hat{r}+\frac{R}{2}|^{2}/\sigma^{2}}}},\quad\hat{\theta}_{2}=\frac{e^{-|\hat{r}+\frac{R}{2}|^{2}/\sigma^{2}}}{e^{-|\hat{r}-\frac{R}{2}|^{2}/\sigma^{2}}+{e^{-|\hat{r}+\frac{R}{2}|^{2}/\sigma^{2}}}}.$ (79)
The final phase-space Hamiltonian is defined above in Eqs. 73-79. To find the eigenvalues of such a Hamiltonian, we construct a joint grid of nuclear positions and momenta $(R_{i},P_{j})$; we include $N_{R}=100$ regular grid points for nuclear coordinate $R\in\left[2,4\right]\textup{\AA}$ and $N_{P}=100$ grid points for nuclear momentum $P\in\left[-\pi N_{P}/\Delta R,...,\pi(N_{P}-1)/\Delta R\right]$ atomic unit (au) based on the conjugate Fourier transform. We then take the following steps in in analogy to the BO theory (but now in phase-space):
  1. Loop over $R_{i}$
    1. Loop over $P_{j}$
      1. Diagonalize the phase-space Hamiltonian (Eq. 77) associated to each joint grid point $\hat{H}_{W}^{\rm PS}(R_{i},P_{j})\ket{\psi_{0}^{\rm PS}}=E^{PS}_{W,0}(R_{i},P_{j})\ket{\psi_{0}^{\rm PS}}$.
    2. End loop over $P_{j}$
    1. Loop over $R_{j}$
      1. Invert the lowest phase-space eigenenergy $E^{PS}_{W,0}(R_{i},P_{j})$ to a function in real space $E_{0}^{\rm PS}(R_{i},R_{j})$ using Eq. 58.
    2. End loop over $R_{j}$
  2. End loop over $R_{i}$
  1. Diagonalize the matrix $E_{0}^{PS}$ (which is a matrix over the nuclear grid in ${\bm{R}}$ with matrix elements $E_{0}^{PS}(R_{i},R_{j})$).
For the parameter $\sigma$ in Eq. 79, we have chosen $\sigma=1.0$ au. That being said, the results are not very sensitive to the value of $\sigma$ in this model problem. For an analysis of the influence of $\sigma$ on our results in this model system, see Appendix A.
5.2.4 Harmonic Results
For the most complete analysis, we will also report results where we consider harmonic BO and phase-space surfaces, i.e. where we force $E_{0}(R)$ or $E^{\rm PS}_{W,0}(R,P)$ to be a quadratic function. For the phase-space surface $E^{\rm PS}_{W,0}(R,P)$, we force the energy to be quadratic with respect to both $P$ and $R$ and expand $E^{\rm PS}_{W,0}$ at the minimum energy point $(R_{\rm min},0)$:
$E^{\rm PS,harmonic}_{W,0}=\frac{\partial^{2}E^{\rm PS}_{W,0}}{\partial P^{2}}\bigg{|}_{P=0}P^{2}+\frac{\partial^{2}E^{\rm PS}_{W,0}}{\partial R^{2}}\bigg{|}_{R=R_{\rm min}}R^{2},$ (80)
where the expansion coefficients are calculated by finite difference. Note that since we have not considered spin-related couplings and there is only one vibrational mode, the phase-space energy will always be minimized at $P=0$ and the off-diagonal Hessian will be 0, i.e., $\frac{\partial^{2}E^{\rm PS}_{W,0}}{\partial R\partial P}=0$. The phase-space harmonic vibrational energy gap is then computed by
$\Delta E=\hbar\sqrt{\frac{\partial^{2}E^{\rm PS}_{W,0}}{\partial R^{2}}\frac{\partial^{2}E^{\rm PS}_{W,0}}{\partial P^{2}}}.$ (81)
5.3 Results
@@
To begin our analysis, in Fig. 1(b), we plot the ground BO potential energy surface that arises from treating the light and heavy degrees of freedom separately, i.e., treating the oxygen atoms as nuclei and the hydrogen atom as an electron. One should not expect this BO separation will be a good approximation if the mass ratio $M/m$ is not large enough, such as the actual masses of oxygen and hydrogen atoms (16 and 1 atomic mass units, respectively). Thus, by varying the nuclei/electron mass ratio $M/m$, we can control non-adiabatic effects for this model Hamiltonian.
@@
In Fig. 2, we report our results for the ground to first excited vibrational energy gap as a function of the different mass ratio. For all methods, the absolute values of the energy gap decreases as the mass ratio increases and converge to the same results in the large mass ratio limit ($\sim$1000 corresponding to a typical nuclear/electronic mass ratio) in Fig. 2(a). Thus, in Fig. 2(b), we plot the error of the energy gap as compared to the exact results. In the small mass ratio limit ($\sim$10 corresponding to the oxygen and hydrogen atoms), one can clearly identify that the phase-space method shows a much smaller error compared to the BO method. Notably, even within a harmonic approximation, the phase-space method still outperforms the BO and sometimes can be comparable to the quantum BO results. Lastly, to best understand these results, the inset of Fig. 2(b) shows the relative errors in the logarithm scale. As one might expect, the phase-space and BO energy errors both scale with the same power law ($\sim m/M$), but the phase-space method consistently gives an order of magnitude smaller relative error even in the large mass ration limit.
@@
It must be emphasized that, with a phase-space Hamiltonian, one can extract other useful properties as well (besides energy) and these quantities are also more accurate than the corresponding BO Hamiltonian. For example, in Fig. 3, we show the errors in the expectation values of a few nuclear and electronic observables. All observables (nuclear or electronic) for both the exact and BO calculations are calculated on a grid using the final wavefunctions following standard procedures. For phase-space calculations, nuclear observables $\langle O\rangle_{mn}$ are obtained by evaluating $\langle O\rangle_{mn}=\bra{\chi_{m}}\hat{O}\ket{\chi_{n}}$ where $\chi_{m}$ and $\chi_{n}$ are the nuclear (vibrational) wavefunctions as determined from the final step of the flowchart in Sec. 5.2.3. As for electronic observables, the expectation value $\langle o\rangle$ is calculated by $(i)$ evaluating the electronic observables on the phase-space grid $(R_{i},P_{j})$ with the ground state phase-space electronic wavefunctions |$\ket{\psi_{0}^{\rm PS}(R_{i},P_{j})}$,
$o_{W}(R_{i},P_{j})=\bra{\psi_{0}^{\rm PS}(R_{i},P_{j})}\hat{o}\ket{\psi_{0}^{\rm PS}(R_{i},P_{j})}$ (82)
and then $(ii)$ averaging over the relevant nuclear Wigner distribution functions (e.g., which are calculated from Eq. 55, $\rho_{mn}^{W}(R,P)=\mathcal{W}\left[\ket{\chi_{m}}\bra{\chi_{n}}\right]$) :
[…]
Let us now evaluate the results.
First, for the electronic momentum, it is important to note that BO calculations always predict zero, as $\bra{\psi_{k}}\hat{\bm{p}}_{\rm e}\ket{\psi_{k}}=0$ for all BO electronic states |$\ket{\psi_{k}}$ at any nuclear geometry. However, as shown in Fig. 3(a), the phase-space method can capture the electronic momentum between the ground and first excited vibrational states with quantitative accuracy. Second, for nuclear momentum, Fig. 3(b) demonstrates that phase-space results are also much more accurate than BO methods. One interesting nuance is that one must replace the nuclear momentum operator $\hat{\bm{P}}$ by $\hat{\bm{P}}-i\hbar\hat{\bm{\Gamma}}$ in the phase-space formalism which represents a canonical – kinetic momentum correspondence. Overall, for the off-diagonal electronic and nuclear momentum (between the same nuclear vibrational eigenstates), the phase-space method performs very well even with the harmonic approximation. As a side note, the on-diagonal nuclear and electronic momentum always vanish due to time reversal symmetry of the whole system (nuclei + electron). Finally, in Fig. 3(c) and (d), while the nuclear and electronic position observables predicted by BO and phase-space methods are not very different, the phase-space methods do provide slightly better results for the excited vibrational states. For further analysis, and a decomposition of the importance of different terms in the phase-space Hamiltonian, see Appendix B.
@@
6 Discussion: Implications for Larger Systems
@@
Of course, one might wonder whether the results here can be generalized to larger systems and be of practical importance. After all, Wigner distributions transformations are notoriously expensive to generate. That being said, however, within the harmonic approximation, there is no reason why the present approach should not be immediately practical for very large systems. After all, whereas standard BO theory diagonalizes the Hessian,
$\mathbf{H}_{\rm BO}=\left(\frac{\partial E_{\rm BO}}{\partial R_{i}\partial R_{j}}\right)$ (84)
within phase-space theory, one must diagonalize (symplectically) a matrix that is only twice as large:
$\mathbf{H}_{\rm PS}=\left(\begin{array}[]{cc}\frac{\partial E_{\rm PS}}{\partial R_{i}\partial R_{j}}&\frac{\partial E_{\rm PS}}{\partial R_{i}\partial P_{j}}\\ \frac{\partial E_{\rm PS}}{\partial P_{i}\partial R_{j}}&\frac{\partial E_{\rm PS}}{\partial P_{i}\partial P_{j}}\end{array}\right)$ (87)
Thus, one can certainly anticipate generating phonons using electronic phase-space Hamiltonians in the immediate future.
In this context, one question that may well arise is how to extract vibrational energies and separate out the center of mass motion, as in Eqs. 80-81 above. For instance, should one work with internal coordinates as in Eq. 77? Alternatively, one could set $P_{\rm NCM}=0$. The latter is size-extensive and likely the most stable. To that end, in Appendix C, we plot results where we do not throw out the first term in Eq. 75, but rather set $P_{\rm NCM}$ to zero and diagonalize the Hamiltonian:
$\hat{H}_{\rm PS}(R,P)=\frac{\left(P-i\hbar\frac{(\hat{\Gamma}_{1}-\hat{\Gamma}_{2})}{2}\right)^{2}}{2\mu}+\frac{\hat{p}^{2}}{2(M+M)}+\frac{\hat{p}^{2}}{2m}+\hat{V}(R,\hat{r}).$ (88)
From the results in Fig. 6, we gather that PS methods still strongly outperforms BO theory – though the results are not quite as strong as they were for Eq. 77. Future work will need to address the pros and cons of these two internal Hamiltonians. As a side note, we mention that our previous work33 with VCD spectroscopy used the full Hamiltonian (rather than the internal Hamiltonian without translation).
Finally, beyond vibrational Hamiltonians, it is worth noting that future work with large systems may well involve excited electronic states, and the present approach should be very useful for so-called phase-space surface hopping dynamics39, 40 where we run classical dynamics on phase-space surfaces, with occasional jumps or hops between surface. To that end, it is worth noting that, for the present model, as shown in Fig. 7 in Appendix D, we find that excited state energies are also improved by using phase-space electronic Hamiltonians. Thus, we anticipate that the present phase-space approach will also be very useful for future nonadiabatic calculations based on classical equations of motion (that do not require a Wigner transform).
7 Future Work and Conclusions
@@
Looking forward, it would appear clear from the plethora of data collected here and in Refs.23, 24 that a phase-space approach is more accurate than a BO approach – and with roughly the same cost. That being said, one must wonder: in what circumstances will the phase-space approach yield meaningfully different answers than BO theory (which performs quite well in most circumstances)? To that end, one can imagine several exciting future problems where one would like to go beyond the limitation of BO theory. The first potential arena would be to study proton transfer41 and proton coupled electron/energy transfer42, 43 where one could combine a phase-space approach with electronic structure methods that treat both light nuclei (protons) and electrons quantum mechanically (e.g., multi-component methods44, 45, 46 and nuclear-electronic orbital methods (NEO)47, 48). Second, it will be fruitful to apply a phase-space approach to study magnetic molecules and solids, where spin-symmetry breaking can lead to intrinsic chiral phonon modes and alter vibrational frequencies49, 50, and electronic momentum and angular momentum are known to be critically important in such circumstances51, 52. Third, given the ab initio and semiclassical nature of the phase-space electronic Hamiltonian, one immediate goal is to investigate nuclear (vibrational) effects at curve crossings during electron transfer processes53 and decipher whether phase-space approaches (that include angular momentum conservation54) can help explain the nature of nuclear effects in chiral-induced spin selectivity55, 56, 57. Fourth, the present phase-space approach opens new and inexpensive avenues for exploring electron-phonon interactions, which could lead to a better atomistic understanding of superconductivity58, 59, 60. At the end of the day, there is a very real chance that phase-space electronic Hamiltonian approaches are here to stay and that a new world of very novel chemistry awaits us.
8 Acknowledgements
This work was supported by the U.S. Air Force Office of Scientific Research (AFOSR) under Grant No. FA9550-23-1-0368 and No. FA9550-18-1-420 (J.E.S.).
Appendix A Influence of the $\sigma$ parameter in our model problem
As shown in Fig. 4, our results are effectively unchanged for small width values, $\sigma=0.5$ au and $\sigma=1.0$ au, which represent the localized $\hat{\Gamma}$ limit; in both cases, these results vastly outperform BO. For large $\sigma$ values (where we do not enforce strong locality), e.g., $\sigma=5.0$ au, the phase-space methods simply recover the BO results.
Figure 4: (a) The ground to first excited vibrational energy gap and (b) the errors in that gap as calculated for BO theory and phase-space methods (i.e. with different $\sigma$ values).
Appendix B Decomposition of the phase-space electronic Hamiltonian
Finally, to gain a better understanding of the phase-space Hamiltonian in vibrational problems, we can decompose the phase-space correction terms in Eq. 40 into two parts: a $\hat{\bm{\Gamma}}\cdot\bm{P}$ contribution and a $\hat{\bm{\Gamma}}\cdot\hat{\bm{\Gamma}}$ contribution. Then, we performed phase-space calculations for which only one of these two terms is included in the Hamiltonian.
Figure 5: Errors of (a) the ground state absolute energy and (b) ground to first excited vibrational energy gap for (1) BO (2), BO+DBOC and (3) phase-space calculations (with different contributions) as compared against exact results.
In Fig. 5(a), we present the errors in the energies for the ground vibrational state ($E_{0}$) as calculated using different methods; to be clear, here we do not consider the error in the energy gap ($E_{1}-E_{0}$) but rather the error of the ground state alone ($E_{0}$). Overall, we find that both (i) the complete phase-space results (using $\hat{\bm{\Gamma}}\cdot\hat{\bm{\Gamma}}$ and $\hat{\bm{\Gamma}}\cdot\bm{P}$ from Eq. 34, purple lines) and (ii) the partial phase-space results using $\hat{\bm{\Gamma}}\cdot\hat{\bm{\Gamma}}$ (triangles) approach the BO plus diagonal Born-Oppenheimer correction (DBOC) results. Obviously, as far as the absolute ground state energy is concerned, the $\hat{\bm{\Gamma}}\cdot\hat{\bm{\Gamma}}$ term is the most important correction.
Next, we turn to the energy gap, $E_{1}-E_{0}$. In Fig. 5(a), we plot the errors in the energy gap for the same methods. Here, we find that neither $\hat{\bm{\Gamma}}\cdot\hat{\bm{\Gamma}}$ nor BO+DBOC offer any improvement. For these problems, we clearly see that including a momentum-dependent potential $\hat{\bm{\Gamma}}\cdot\bm{P}$ is critical for recovering an accurate energy gap. In a sense, these results should be intuitive. After all, in the hydrogen atom case, $\hat{\bm{\Gamma}}\cdot\bm{P}$ and $\hat{\bm{\Gamma}}\cdot\hat{\bm{\Gamma}}$ correspond exactly to adiabatic (i.e., DBOC) and non-adiabatic corrections to BO theory as discussed in Refs. 9, 10. Furthermore, similar results as far the effect of the DBOC on absolute and relative energies were reported by Hammes-Schiffer et al.61
Appendix C Results without eliminating the MCM motion
We have performed the same calculation using a phase-space Hamiltonian (Eq. 88) that does not eliminate the MCM motion but instead imposes the condition $P_{\rm NCM}=0$. Both phase-space results, calculated from Eq. 77 and Eq. 88, outperform the BO results, with the removal of MCM motion yielding better accuracy. For large systems, with large total mass $M$, these two methods should be very similar. That being said, formally speaking, Eq. 40 is size consistent (meaning that the vibrational energies for a molecule will be the same if we one or many noninteracting molecules)– but this size consistency is eliminated if we remove the total translational energy as in Eq. 88 and set $\tilde{H}_{PS}=\hat{H}_{PS}-\frac{\hat{p}^{2}}{2\sum_{A}M_{A}}$. For this reason, the results using Eq. 77 are more likely to be relevant when investigating large systems.
Figure 6: (a) The ground to first excited vibrational energy gap and (b) the errors in that gap as calculated for BO theory and phase-space methods (pink circle with Eq. 77 and purple box with Eq. 88). Both phase space approaches strongly outperform BO.
Appendix D Vibrational energies calculated for both ground and excited electronic surfaces
We have also calculated the errors in the vibrational energy gaps for a mass ratio of $M/m=16$ in Eq. 77. The phase-space methods show improvement in the vibrational energies on both the ground and excited electronic energy surfaces, highlighting another area where phase-space approaches may well find application.
Figure 7: The errors in the vibrational energy gaps $\Delta E_{N}=E_{N}-E_{0}$ are calculated for both BO theory and phase-space methods. The vibrational states are ordered by their energies, with $E_{N}$ representing the energy of the $N$-th vibrational states irrespective of whether the latter corresponds to a ground or excited electronic surface. Vibrational states on the ground BO electronic surface are labeled in blue (BO) and green (PS), while those on the excited electronic surface are labeled in red (BO) and yellow (PS). Note that phase space approaches outperform BO both for ground and excited state vibrational energies.